OpenStudy (anonymous):

What is square root of 150 in simplified radical form? A. Irrational Square Roots B. Irrational Square Roots C. Irrational Square Roots D. Irrational Square Roots

OpenStudy (anonymous):

sqaure root of 150 meaning \[\sqrt{150}\]

OpenStudy (anonymous):

@Smita12

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Break it into factors.

OpenStudy (igreen):

6 and 25

OpenStudy (igreen):

6 * sqrt 25

OpenStudy (igreen):

What's the square root of 25?

OpenStudy (anonymous):

oh it would be \[\sqrt[5]{6}\]?

OpenStudy (igreen):

The other way around.

OpenStudy (igreen):

\(6 \times \sqrt{25}\) Simplify the square root of 25

OpenStudy (anonymous):

\[\sqrt[25]{6}\]

OpenStudy (igreen):

No.. what's the square root of 25?

OpenStudy (anonymous):

5

OpenStudy (igreen):

Oh, wait that's: \(\sqrt{6} \times \sqrt{25}\) Simplifying 25 gives us 5, like you said: \(\sqrt{6} \times 5\) or \(5 \sqrt6\)

OpenStudy (anonymous):

Thank you ^_^

OpenStudy (igreen):

No problem.

OpenStudy (anonymous):

my teacher looked at it online and said i got some wrong but not all of them and said i would most likely get a 80% or higher which i took the test fast and got 80%

OpenStudy (anonymous):

it was a study guide that you helpped me with